# A spring–block theory of feature learning in deep neural networks

### v2

@@ Abstract
Feature-learning deep nets progressively collapse data to a regular low-dimensional geometry. How this phenomenon emerges from collective action of nonlinearity, noise, learning rate, and other choices that shape the dynamics, has eluded first-principles theories built from microscopic neuronal dynamics. We exhibit a noise–nonlinearity phase diagram that identifies regimes where shallow or deep layers learn more effectively. We then propose a macroscopic mechanical theory that reproduces the diagram, explaining why some DNNs are lazy and some active, and linking feature learning across layers to generalization.
@@
Deep neural networks (DNNs) progressively compute features from which the final layer generates predictions. Something remarkable happens when they are optimized via stochastic dynamics over a data-dependent energy: Each layer learns to compute better features than the previous one [1], ultimately transforming the data to a regular low-dimensional geometry [2, 3, 4, 5, 6, 7]. Feature learning is a striking departure from kernel machines or random feature models (RFM) which compute linear functions of a fixed set of features [8, 9, 10]. How it emerges from microscopic interactions between millions or billions of artificial neurons remains a central open question in deep learning [11, 12, 13, 14, 15].
@@
Even with a single hidden layer [16, 17, 18, 19], the interplay between initialization [20], width [21, 22, 23], learning rate [24, 25], batch size [26, 27], and data [28, 29, 30] results in a bewildering range of training dynamics. Deeper networks can be analyzed in asymptotic regimes where they resemble kernel machines [31, 32], with simplified layer-wise training [33, 34], and by removing non-linear activations [35, 36]. There have been exciting recent advances on deep mean field theory in the infinite-width limit [37, 38] and on effective low-dimensional SGD dynamics  [39, 40]. These are tremendously useful to build insight, but the full deep, non-linear setting has thus far challenged a statistical mechanics approach where feature learning would emerge from microscopic interactions. There is thus a gap between current theory and the DNNs used in practice and a change of perspective may be useful.
@@
In this paper we rather take a thermodynamical, top-down approach and look for a simplest well-understood phenomenological model which captures the essential aspects of feature learning. We first show that DNNs can be mapped to a phase diagram defined by noise and nonlinearity, with phases where layers learn features at equal rates, and where deep or shallow layers learn better features. “Better” is quantified through the notion of data separation—the ratio of feature variance within and across classes. To explain this phase diagram, we propose a macroscopic theory of feature learning in deep, nonlinear neural networks: we show that the stochastic dynamics of a nonlinear spring–block chain with asymmetric friction fully reproduce the phenomenology of data separation over training epochs and layers. The phase diagram is universal with respect to the source of stochasticity: varying dropout rate, batch size, label noise, or learning rate all result in the same phenomenology.
@@
These findings considerably generalize recent work which showed that in many DNNs each layer improves data separation by the same factor, with surprising regularity [3]. This law of data separation can be proved for linear DNNs with a particular choice of data and initialization [41, 42], but it is puzzling why just as many networks do not abide by it. Fig. 4 illustrates the challenge: the same DNN trained with three different parameter sets results in strikingly different distributions of data separation over layers. Understanding how these non-uniform distributions arise and how it affects generalization is key to understanding feature learning.
Even linear DNNs induce a complex energy landscape and highly nonlinear training dynamics [36] that can result in non-even separation. The relationship between the energy landscape and learning has been studied with tools from statistical mechanics [43, 44, 45, 46], random matrix theory [47], and optimal transport [48, 17] but it remains unclear how to connect these insights to macroscopic feature learning in real DNNs.
@@
Figure 1: Phase diagrams of DNN load curves (red) for nonlinearity vs. (a) data noise, (b) learning rate, (c) dropout, and (d) batch size. In all cases, noise is strongest on the left, and nonlinearity strongest at the top. Background shading encodes test accuracy. Results are averaged over 10 independent runs on MNIST.
@@
In our theory, spring elongations model data separation. The empirical risk exerts a load on the network to which its layers respond by separating the data, subject to nonlinearity modeled by friction. Difference in length between consecutive springs gives a load on the incident block. Asymmetric friction models the propagation of signal and noise in forward and backward passes. It models dynamical nonlinearity in a network which absorbs load (or spoils signal in gradient) causing the shallow layers to “extend less” (do not learn features). Stochasticity, be it from stochastic gradient descent (SGD) [49, 50], dropout [51], or noisy data [52], reequilibrates the load.
@@
The resulting model reproduces the dynamics and the phase diagram of feature learning surprisingly well. It explains when data separation is uniformly distributed across layers but also when deep or shallow layers learn faster. It shows why depth may be problematic for learning and why nonlinearity is a double-edged sword, resulting in expressive models but facilitating overfitting. A stability argument suggests a link between generalization and layerwise data separation which we remarkably find true in real DNNs.
@@
for $\ell=1,\ldots,L$. The activation-dependent normalization factors $c_{\ell}$ scale the variance of hidden features in each layer close to $1$ [53, 54, 55]; they can be replaced by batch normalization [56].
It is natural to expect that in a well-trained DNN the intermediate features $\bm{{x}}_{\ell}$ improve progressively over $\ell$. Following recent work on neural collapse we measure separation as the ratio of variance within and across classes [2, 3, 4, 5, 41]; in supplemental material (SM) [57] we show that analogous phenomenology exists in regression. Let $\mathcal{X}_{\ell}^{k}$ collect the $\ell$th postactivation for examples from class $k$, with $\bar{\bm{{x}}}_{\ell}^{k}$ and $N_{k}$ its mean and cardinality. The between-class and within-class covariances are
$\Sigma_{\ell}^{\mathrm{w}}:=\operatorname*{Ave}_{k}\,(\mathrm{Cov}\mathcal{X}_{\ell}^{k}),\quad\quad\Sigma_{\ell}^{\mathrm{b}}:=\mathrm{Cov}\left(\tfrac{N_{k}}{N}\bar{\bm{{x}}}_{\ell}^{k}\right)_{k=1}^{K},$ (2)
where $\mathrm{Ave}$ is the average; $\Sigma^{\mathrm{b}}_{\ell}$ is the between-class “signal” for classification, and $\Sigma^{\mathrm{w}}_{\ell}$ is the within-class variability. Data separation at layer $\ell$ is then defined as
@@
The difference $d_{\ell}=D_{\ell-1}-D_{\ell}$ represents the contribution of the $\ell$th layer. We call the “discrete curve” $D_{\ell}$ vs. $\ell$ the load curve in anticipation of the mechanical analogy. If indeed each layer improves the representation the load curve should monotonically decrease.
Modern overparameterized DNNs may perfectly fit the training data for different choices of hyperparameters, while yielding different load curves. An extreme is an RFM where only the last layer learns while the rest are frozen at initialization [9, 58]. As the entire load is concentrated on one layer, one might intuitively expect that a more even distribution will result in better performance.
@@
which results in a linear load curve [59]. This can be proved in linear DNNs with orthogonal initialization and gradient flow training [60, 41, 42], but as we show below it is a brittle phenomenon: nonlinear activations immediately break the balance. Equiseparation thus requires additional ingredients; He and Su highlight the importance of an appropriate learning rate.
The nonlinearity–noise phase diagram—
We show that DNNs define a family of phase diagrams such that (i) increasing nonlinearity results in increasingly “concave” load curves, with deeper layers learning better features and thus taking a higher load, $d_{\ell}<d_{\ell^{\prime}}$ for $\ell<\ell^{\prime}$; (ii) adding noise to the dynamics rebalances the load across layers; and (iii) further increasing noise results in convex load curves, with shallower layers learning better features.
We report these findings in Fig. 1. In all panels the abscissa measures stochasticity while the ordinate measures nonlinearity. To measure nonlinearity we use the slope $\alpha\in[0,1]$ of the negative part of a LeakyReLU activation, $\mathrm{LReLU}(x):=\max(\alpha x,x)$; $\alpha=1$ gives a linear DNN, $\alpha=0$ the ReLU. Consider for example Fig. 1(a) where noise is introduced in labels and data (we randomly reset a fraction $p$ of the labels $\mathbf{y}$, and add Gaussian noise with variance $p^{2}$ to $\bm{{x}}_{0}$). Without noise (upper right corner), the load curve is concave; this resembles an RFM or a kernel machine. Increasing noise (right to left) yields a linear and then a convex load curve.
Interestingly, the same phenomenology results from varying learning rate, dropout, and batch size (Fig. 1(b, c, d)). Lower learning rate, smaller dropout, and larger batch size all indicate less stochasticity in dynamics. Martin and Mahoney [61] call them temperature-like parameters in the statistical mechanics sense; see also Zhou et al. [62]. In all cases high nonlinearity and low noise result in “lazy” training [63] and a concave load curve. Low nonlinearity and high noise (bottom-left) result in convex load curves. We observe the same behavior for other datasets (e.g., CIFAR10) and network architectures (e.g., CNNs); cf. SM [57]. This phenomenology is at the heart of feature learning. It asks how we go from non-feature learning models like RFM or NTK (where only or mostly deep layers learn) to models where all layers learn.
@@
We now show something surprising: the complete phenomenology of feature learning, as seen through layerwise data separation, can be reproduced by the stochastic dynamics of a simple spring–block chain. As in Fig. 2, we interpret $d_{\ell}$, the signed elongation of the $\ell$th Hookean spring, as data separation afforded by the $\ell$th layer. Block movement is impeded by friction which models the effect of dynamical nonlinearity on gradients; noise in the force models stochasticity from mini-batching, dropout, or elsewhere.
@@
for $\ell=1,\ldots,L$, where we assumed unit masses, $k$ is the elastic modulus, $\gamma$ a linear damping coefficient sufficiently large to avoid oscillations, and $\xi_{\ell}$ is noise such that $\langle\xi_{\ell}(t)\xi_{\ell}(s)\rangle=\delta(t-s)$. As DNNs at initialization do not separate data we let $x_{\ell}(0)=0$. The dynamics is driven by force applied to the last block $F=k(y-x_{L})$. We set $x_{0}\equiv 0$ and $x_{L+1}\equiv y$ to model training data and targets. The load curve plots the distance of the $\ell$th block from the target $D_{\ell}=y-x_{\ell}$. It reflects how much the $\ell$th spring—or the $\ell$th layer—contributes to “explaining” the total extension—or the total data separation. One key insight is that the friction must be asymmetric to model the propagation of noise during training as we elaborate below. We set the sliding and maximum static friction to $\mu_{\rightarrow}$ for rightward and $\mu_{\leftarrow}$ for leftward movement. In this model the friction acts on noisy force. If noise is added to the velocity independently of the friction we obtain a more standard Langevin dynamics. Although this is physically less realistic it shows similar qualitative behavior (for additional details see SM [57]).
@@
We now show experimentally and analytically that the proposed model results in a phenomenology analogous to that of data separation in real networks. Indeed, there is a striking similarity between the phase diagram of the spring–block model in Fig. 3 and the DNN phase diagrams in Fig. 1. Surprisingly, it is not only the equilibria of the two systems that are similar, but also the stochastic dynamics; we show this in Fig. 4.
@@
1: nonlinearity breaks the separation balance.
Let us first analyze how our model explains concave load curves and lazy training. For simplicity we work in the overdamped approximation $\gamma\gg 1$; in SM [57] we show that the second order system has the same qualitative behavior. Scaling time by $\gamma$, Eq. (4) then yields
@@
where $(\bm{L}x)_{\ell}:=x_{\ell+1}-2x_{\ell}+x_{\ell-1}$ and
[…]
@@
if we assume the initial elastic potential $k(y-x_{0})^{2}/2$ is sufficiently large such that all blocks will eventually move. In this case $\Delta d^{*}_{\ell}:=d^{*}_{\ell+1}-d^{*}_{\ell}=(\bm{{L}}x^{*})_{\ell}={\mu_{\rightarrow}}/{k}>0$ and the load curve is concave. Note that this result only involves $\mu_{\rightarrow}$ but not $\mu_{\leftarrow}$ as without noise and with sufficient damping the blocks only move to the right. The interpretation is that in equilibrium, the friction at each block absorbs some of the load so that the shallower springs extend less; in a DNN this corresponds to lazy training where the deepest layer, whose gradients do not experience nonlinearity, does most of the separation.
Figure 4: The load curves at convergence (a) and trajectories (b, c, d) for an $7-$hidden layer ReLU fully connected DNN on MNIST (left) and the proposed spring–block model (right). For the DNN (left), the ordinate is $D_{\ell}$ (data separation at layer $\ell$); the dashed line is the training loss. Characteristic behaviors: (b) high nonlinearity (high friction) and low randomness in training (noise in force); (c): balanced nonlinearity (friction) and randomness (force noise); (d) low nonlinearity (friction) and large training randomness (force noise). In the spring system (right), the ordinate represents the distance to the target $D_{\ell}=y-x_{\ell}$ for colored lines. These values are scaled to match the same regime used in the DNN for visualization purposes. The dashed line represents the force at the rightmost block $F$ which is at a different scale from $D_{\ell}$.
2: Noise reequilibrates the load.
If friction reduces load, how can then a chain with friction—or a nonlinear DNN—result in a uniformly distributed load? We know from Fig. 1 that in DNNs stochastic training helps achieve this. We now show how our model reproduces this behavior and in particular a counterintuitive phenomenon in Fig. 1 where large noise results in convex load curves where shallow layers learn better than deep. We show that this happens when $\mu_{\leftarrow}>\mu_{\rightarrow}$.
@@
Further, when $\Delta d\in[\frac{\mu_{\rightarrow}-\mu_{\leftarrow}}{2k},\frac{\mu_{\rightarrow}}{k}]$, it holds that $\frac{\mathrm{d}\mu_{\text{eff}}}{\mathrm{d}\varepsilon}\leq 0$. It implies that increasing noise always decreases effective friction which resembles phenomena like acoustic lubrication or noise-induced superlubricity [64]. When $\mu_{\leftarrow}>\mu_{\rightarrow}>0$, as we vary $\varepsilon$ from 00 to $\infty$ we will first see a concave, then a linear, and finally a convex load curve, corresponding to the transition from lazy to active training in DNNs [63]. As a result, when $\mu_{\leftarrow}>\mu_{\rightarrow}>0$ our model explains the entire DNN phase diagram.
How can we relate the condition $\mu_{\leftarrow}>\mu_{\rightarrow}>0$ to DNN phenomenology? Note that in our model $\mu_{\leftarrow}$ is activated only due to the noise, since the signal (the pulling force) is always to the right. In a DNN, the forward pass of the backpropagation algorithm computes the activations while the backward pass computes the gradients and updates the weights. In the forward pass, the input is multiplied by the weight matrices starting from the shallowest to the deepest. With noise (e.g. dropout), the activations of the deepest layer accumulate the largest noise. Thus noise in DNN training chiefly propagates from shallow to deep. This is exactly what $\mu_{\leftarrow}>\mu_{\rightarrow}$ in our model implies, since $\mu_{\leftarrow}$ is only triggered by noise.
Indeed, from Eq. (8) we see that when only friction for leftward move is present, $\mu_{\leftarrow}>\mu_{\rightarrow}=0$, the load curve is convex if $\varepsilon>0$ and linear if $\varepsilon=0$ (Fig. 5(b)). This equilibrium is independent of $\mu_{\leftarrow}$ since no block moves left. This observation parallels findings in linear DNNs where training with gradient flow and “whitened” data leads to a linear load curve [41]. In contrast, introducing dropout or large learning rate makes the load curve convex (Fig. 5(a)). This shows that dynamical nonlinearity—or friction in our model—exists even in linear DNNs, so that their learning dynamics are nonlinear [36].
Figure 5: (a) Load curve of a deep linear neural network on random orthonormal datasets. (b) Spring–block model without left friction ($\mu_{\rightarrow}=0$) with $\mu_{\leftarrow}=0.1$, under varying levels of noise $\varepsilon$. (c) An illustration of asymmetric friction.
@@
We finally give an example of how the proposed theory yields practical insights. Among all spring–block chains under a fixed load, the equiseparated one is the most stable in the sense of having the lowest potential energy. This motivates us to look at the test accuracy of DNNs as a function of load curve curvature; this is shown as background shading in Fig. 1. The result is intriguing: Linear load curves correspond to the highest test accuracy. It suggests that by balancing nonlinearity with noise, DNNs are at once highly expressive and not overfitting. That a uniform load distribution yields the best performance is intuitively pleasing. It is also valuable operationally as it may help guide training to find better networks. We show an example involving convolutional networks in SM [57].

### v4

# Spring–block theory of feature learning in deep neural networks

@@ Abstract
Feature-learning deep nets progressively collapse data to a regular low-dimensional geometry. How this emerges from the collective action of nonlinearity, noise, learning rate, and other factors, has eluded first-principles theories built from microscopic neuronal dynamics. We exhibit a noise–nonlinearity phase diagram that identifies regimes where shallow or deep layers learn more effectively and propose a macroscopic mechanical theory that reproduces the diagram and links feature learning across layers to generalization.
@@
Deep neural networks (DNNs) progressively compute features from which the final layer generates predictions. When optimized via stochastic dynamics over a data-dependent energy, each layer learns to compute better features than the previous one [1], ultimately transforming the data to a regular low-dimensional geometry [2, 3, 4, 5, 6, 7]. Feature learning is a striking departure from kernel machines or random feature models (RFM) which compute linear functions of fixed features [8, 9, 10]. How it emerges from microscopic interactions between millions of artificial neurons is a central open question in deep learning [11, 12, 13, 14, 15].
@@
Even with a single hidden layer [16, 17, 18, 19], the interplay between initialization [20], width [21, 22, 23], learning rate [24, 25], batch size [26, 27], and data [28, 29, 30] results in a bewildering range of training dynamics. Deeper networks can be analyzed in various asymptotic regimes [31, 32], with simplified training [33, 34], or without non-linearity [35, 36]. There have been exciting advances in the infinite-width limit [37, 38] and on low-dimensional SGD dynamics [39, 40]. These give invaluable insight, but the full deep, non-linear setting has eluded a statistical mechanics approach where features emerge from microscopic interactions. A change of perspective may help close this gap.
@@
In this paper we take a thermodynamical, top-down approach and look for a simple phenomenological model which captures the feature learning phenomenology. We show that DNNs can be mapped to a phase diagram defined by noise and nonlinearity, with phases where layers learn features uniformly, and where deep or shallow layers learn better. “Better” is quantified through data separation—the ratio of feature variance within and across classes. To explain this phase diagram, we propose a macroscopic theory of feature learning in deep, nonlinear neural networks: we show that the stochastic dynamics of a nonlinear spring–block chain with asymmetric friction fully reproduce the phenomenology of data separation over training epochs and layers. The phase diagram is universal with respect to the source of stochasticity: varying dropout rate, batch size, label noise, or learning rate all result in the same phenomenology.
@@
Our findings generalize recent work showing that in many DNNs each layer improves data separation by the same factor, with surprising regularity [3]. This law of data separation can be proved for linear DNNs with a particular choice of data and initialization [41, 42]. It is however puzzling why just as many networks do not abide by it. FIG. 4 shows that training the same DNN with three different parameter sets results in strikingly different distributions of data separation over layers. Even linear DNNs induce a complex energy landscape [43, 44, 45, 46, 47, 48, 17] and nonlinear training dynamics [36] that can result in non-even separation. Understanding why this happens and how it affects generalization is key to understanding feature learning.
@@
Figure 1: Phase diagrams of DNN training load curves (red) for nonlinearity vs. (a) data noise, (b) learning rate, (c) dropout, and (d) batch size. The non-linearity is controlled by the negative slope in LeakyReLU, with values of 1, 0.8, 0.6, 0.4, 0.2, and 0 from the bottom row to the top row. In all cases, noise is strongest on the left, and nonlinearity strongest at the top. Background shading encodes test accuracy. Results are averaged over 10 independent runs on MNIST.
@@
In our theory, spring elongations model data separation. The empirical risk exerts a load on the network to which its layers respond by separating the data, subject to nonlinearity modeled by friction. Difference in length between consecutive springs results in a load on the incident block. Friction models dynamical nonlinearity which absorbs load (or spoils the signal in gradients), causing the shallow layers to “extend” (separate data) less. Stochasticity from stochastic gradient descent (SGD) [49, 50], dropout [51], or noisy data [52], reequilibrates the load.
@@
The resulting model reproduces the dynamics and the phase diagram of feature learning surprisingly well. It explains when data separation is uniformly distributed across layers and when deep or shallow layers learn faster. It shows why depth may hurt and why nonlinearity is a double-edged sword, resulting in expressive models but facilitating overfitting. A stability argument suggests a link between generalization and layerwise data separation which we remarkably find enacted in real DNNs. This observation is of great practical interest: together with our understanding of noise and nonlinearity it suggests a simple strategy for hyperparameter tuning and model selection which may be a compelling alternative to grid-search approaches; we explore this in FIG. 6 and in further experiments in SM [53].
@@
for $\ell=1,\ldots,L$. We call the layers with small $\ell$ (near the input) shallow and those with large $L$ (near the output) deep. The activation-dependent normalization factors $c_{\ell}$ scale the variance of hidden features in each layer close to $1$ [54, 55, 56]; they can be replaced by batch normalization [57].
It is natural to expect that in a well-trained DNN the intermediate features $\bm{{x}}_{\ell}$ improve progressively over $\ell$. Following recent work on neural collapse we measure separation as the ratio of variance within and across classes [2, 3, 4, 5, 41]; in supplemental material (SM) [53] we show that analogous phenomenology exists in regression. Let $\mathcal{X}_{\ell}^{k}$ collect the $\ell$th postactivation for examples from class $k$, with $\bar{\bm{{x}}}_{\ell}^{k}$ and $N_{k}$ its mean and cardinality. The within-class and between-class covariances are
$\Sigma_{\ell}^{\mathrm{w}}:=\operatorname*{Ave}_{n_{k}}\,(\mathrm{Cov}\mathcal{X}_{\ell}^{k}),\quad\Sigma_{\ell}^{\mathrm{b}}:=\operatorname*{Cov}_{n_{k}}\left(\bar{\bm{{x}}}_{\ell}^{k}\right)_{k=1}^{K},$ (2)
where $\operatorname*{Ave}_{n_{k}}$ and $\operatorname*{Cov}_{n_{k}}$ are the weighted average and covariance with weight $n_{k}=N_{K}/N$. $\Sigma^{\mathrm{b}}_{\ell}$ is the between-class “signal” for classification, and $\Sigma^{\mathrm{w}}_{\ell}$ is the within-class variability. Data separation at layer $\ell$ is then defined as
@@
The difference $d_{\ell}=D_{\ell-1}-D_{\ell}$ represents the contribution of the $\ell$th layer. We call the “discrete curve” $D_{\ell}$ vs. $\ell$ the load curve in anticipation of the mechanical analogy. If indeed each layer improves the data representation the load curve should monotonically decrease.
Modern overparameterized DNNs may perfectly fit the training data for different choices of hyperparameters, while yielding different load curves. One extreme is an RFM where only the last layer learns while the rest are frozen at initialization [9, 58]. As the entire load is concentrated on one layer, one might intuitively expect that a more even distribution results in better performance.
@@
giving a linear load curve [59]. This can be proved in linear DNNs with orthogonal initialization and gradient flow training [60, 41, 42], but as we show below it is brittle: nonlinearity breaks the balance. Equiseparation thus requires additional ingredients; He and Su highlight the importance of an appropriate learning rate.
The noise–nonlinearity phase diagram—
We show that DNNs define a family of phase diagrams such that (i) increasing nonlinearity results in increasingly “concave” load curves, with deeper layers learning better features (taking a higher load), $d_{\ell}<d_{\ell^{\prime}}$ for $\ell<\ell^{\prime}$; (ii) noise in the dynamics rebalances the load; and (iii) increasing noise results in convex load curves, with shallower layers learning better features.
We report these findings in FIG. 1. In all panels the abscissas measure stochasticity and the ordinates nonlinearity. We control nonlinearity by varying the slope $\alpha\in[0,1]$ of the negative part of a LeakyReLU activation, $\mathrm{LReLU}(x):=\max(\alpha x,x)$; $\alpha=1$ gives a linear DNN, $\alpha=0$ the ReLU. Consider for example FIG. 1(a) where noise is introduced in labels and data (we randomly reset a fraction $p$ of the labels $\mathbf{y}$, and add Gaussian noise with variance $p^{2}$ to $\bm{{x}}_{0}$). Without noise (upper right corner), the load curve is concave; this resembles an RFM or a kernel machine. Increasing noise (right to left) yields a linear and then a convex load curve.
The same phenomenology results from varying learning rate, dropout, and batch size (FIG. 1(b, c, d)). Lower learning rate, smaller dropout, and larger batch size all indicate less stochasticity in dynamics. Martin and Mahoney [61] call them temperature-like parameters; see also Zhou et al. [62]. In SM we give simple rules of thumb for trading off these different sources of noise [53]. In all cases high nonlinearity and low noise result in a concave load curve where parameters of deep layers move faster than those of shallow. Low nonlinearity and high noise (bottom-left) result in convex load curves. We observe the same behavior for other datasets (e.g., CIFAR10) and network architectures (e.g., CNNs); cf. SM [53]. The goal is to understand how varying these parameters transitions from non-feature learning models like RFM or NTK to models where all layers learn.
@@
We now show that the complete phenomenology of feature learning, as seen through layerwise data separation, is mirrored by the stochastic dynamics of a simple spring–block chain. As in FIG. 2, we interpret $d_{\ell}$, the signed elongation of the $\ell$th Hookean spring, as data separation by the $\ell$th layer. Block movement is impeded by friction which models the effect of dynamical nonlinearity on gradients; noise in the force models stochasticity from mini-batching, dropout, or elsewhere.
@@
for $\ell=1,\ldots,L$, where we assumed unit masses, $k$ is the elastic modulus, $\gamma$ is a linear damping coefficient sufficiently large to avoid oscillations, $\epsilon$ controls the noise strength and $\xi_{\ell}$ is noise such that $\langle\xi_{\ell}(t){,}\xi_{\ell}(s)\rangle=\delta(t-s)$. As DNNs at initialization do not separate data we let $x_{\ell}(0)=0$. The dynamics is driven by force applied to the last block $F=k(y-x_{L})$. We set $x_{0}\equiv 0$ and $x_{L+1}\equiv y$ to model training data and targets. The load curve plots the distance of the $\ell$th block from the target $D_{\ell}=y-x_{\ell}$. It reflects how much the $\ell$th spring—or the $\ell$th layer—contributes to “explaining” the total extension—or the total data separation. One key insight is that the friction must be asymmetric to model the propagation of noise during training as we elaborate below. We set the sliding and maximum static friction to $\mu_{\rightarrow}$ for rightward and $\mu_{\leftarrow}$ for leftward movement. In this model the friction acts on noisy force. If noise is added to the velocity independently of the friction we obtain a more standard (but physically less realistic) Langevin dynamics which exhibits similar qualitative behavior ; for additional details see SM [53].
@@
We now show experimentally and analytically that the proposed model results in a phenomenology analogous to that of data separation in real networks. There is a striking similarity between the phase diagram of the spring–block model in FIG. 3 and the DNN phase diagrams in FIG. 1. Not only are the equilibria of the two systems similar, but also the stochastic dynamics; we show this in FIG. 4.
@@
1: Nonlinearity breaks the separation balance—
We first show how our model explains concave load curves. For simplicity we work in the overdamped approximation $\gamma\gg 1$; in SM [53] we show that the second order system has the same qualitative behavior. Scaling time by $\gamma$, Eq. (4) yields
@@
where $(\bm{L}x)_{\ell}:=x_{\ell+1}-2x_{\ell}+x_{\ell-1}$ and $\sigma(z)=0$ if $-\mu_{\leftarrow}\leq z\leq\mu_{\rightarrow}$; $z-\mu_{\rightarrow}$ if $z>\mu_{\rightarrow}$; and $z+\mu_{\leftarrow}$ if $z<-\mu_{\leftarrow}$.
@@
if we assume the initial elastic potential $k(y-x_{0})^{2}/2$ is sufficiently large such that all blocks eventually move. In this case $\Delta d^{*}_{\ell}:=d^{*}_{\ell+1}-d^{*}_{\ell}=(\bm{{L}}x^{*})_{\ell}={\mu_{\rightarrow}}/{k}>0$ and the load curve is concave. Note that this result only involves $\mu_{\rightarrow}$ but not $\mu_{\leftarrow}$ as without noise and with sufficient damping the blocks only move to the right. The interpretation is that in equilibrium, the friction at each block absorbs some of the load so that the shallower springs extend less; in a DNN this corresponds to a regime where the deepest layer, whose gradients do not experience nonlinearity, does most of the separation.
Figure 4: The load curves at convergence (a) and trajectories (b, c, d) for a $7-$hidden layer ReLU MLP on MNIST ( 1) vs our spring–block model ( 2). For the MLP ( 1), the ordinate is $D_{\ell}$ (data separation at layer $\ell$); the dashed line is the training loss. Characteristic behaviors: (b) high nonlinearity (high friction) and low randomness in training (noise in force); (c): balanced nonlinearity and randomness; (d) low nonlinearity and strong randomness. In the spring system ( 2), the ordinate is the distance to the target $D_{\ell}=y-x_{\ell}$. The values are scaled to match the same regime used in the DNN for ease of visualization. The dashed line is the force at the rightmost block $F$ which is at a different scale from $D_{\ell}$.
2: Noise reequilibrates the load—
If friction reduces load, how can a chain with friction—or a nonlinear DNN—result in a uniformly distributed load? We know from FIG. 1 that in DNNs stochastic training helps achieve this. We now show how our model reproduces this behavior and in particular a counterintuitive phenomenon in FIG. 1 where large noise results in convex load curves where shallow layers learn better than deep. We show that this happens when $\mu_{\leftarrow}>\mu_{\rightarrow}$.
@@
Further, when $\Delta d\in[\frac{\mu_{\rightarrow}-\mu_{\leftarrow}}{2k},\frac{\mu_{\rightarrow}}{k}]$, it holds that $\frac{\mathrm{d}\mu_{\text{eff}}}{\mathrm{d}\varepsilon}\leq 0$. It implies that increasing noise always decreases effective friction. This resembles phenomena like acoustic lubrication or noise-induced superlubricity [63]. When $\mu_{\leftarrow}>\mu_{\rightarrow}>0$, as we vary $\varepsilon$ from 00 to $\infty$ we will first see a concave, then a linear, and finally a convex load curve. Therefore, when $\mu_{\leftarrow}>\mu_{\rightarrow}>0$ our model explains the entire DNN phase diagram.
How can we relate the condition $\mu_{\leftarrow}>\mu_{\rightarrow}>0$ to DNN phenomenology? Note that in our model $\mu_{\leftarrow}$ is activated only due to the noise, since the signal (the pulling force) is always to the right. In a DNN, the forward pass of the backpropagation algorithm computes the activations while the backward pass computes the gradients. In the forward pass, the input is multiplied by the weight matrices starting from the shallowest to the deepest. With noise (e.g., dropout), the activations of the deepest layer accumulate the largest noise. Thus noise in DNN training chiefly propagates from shallow to deep, yielding noisiest gradients in deepest layers. This is exactly what $\mu_{\leftarrow}>\mu_{\rightarrow}$ in our model implies, since $\mu_{\leftarrow}$ is only triggered by noise.
Indeed, from Eq. (7) we see that when only friction for leftward movement is present, $\mu_{\leftarrow}>\mu_{\rightarrow}=0$, the load curve is convex if $\varepsilon>0$ and linear if $\varepsilon=0$ (FIG. 5(b)). The equilibrium is independent of $\mu_{\leftarrow}$ since no block moves left. This parallels findings in linear DNNs where training with gradient flow and “whitened” data leads to a linear load curve [41]. In contrast, introducing noise makes the load curve convex (FIG. 5(a)). This shows that dynamical nonlinearity—or friction in our model—exists even in linear DNNs, so that their learning dynamics are nonlinear [36].
Figure 5: (a) Load curve of a deep linear neural network on random orthonormal datasets. (b) Spring–block model without left friction ($\mu_{\rightarrow}=0$) with $\mu_{\leftarrow}=0.1$, under varying levels of noise $\varepsilon$. (c) Load curve of 7-layer CNN trained on the CIFAR10 dataset; dp and ACC denote the dropout ratio and test accuracy. (d) An illustration of asymmetric friction.
@@
We finally show how our theory gives insight into generalization. Among all spring–block chains under a fixed load, the equiseparated one is the most stable in the sense of having the lowest potential energy. This motivates us to study the test accuracy of DNNs—shown as background shading in FIG. 1—as a function of load curve curvature. The result is intriguing: linear load curves correspond to the highest test accuracy. It suggests that by balancing nonlinearity with noise, DNNs are at once highly expressive and not overfitting. That a uniform load distribution yields the best performance is intuitively pleasing, but it is also valuable operationally as it may help guide training to find better networks. In FIG. 6 we show a proof-of-concept example with a CNN: initial training yields a concave load curve. Our theory suggests that more noise in training would flatten the load curve and improve generalization—which is indeed what happens. Further details and experiments with very deep networks can be found in SM [53].